OpenStudy (anonymous):

if a fountains velocity is 240ft per second what is it in miles per min?

OpenStudy (anonymous):

Pleaseeeee!!!

OpenStudy (anonymous):

Conversion factors... one sec

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

\[240 \frac{\cancel{ft}}{\cancel s} * (\frac{60 \cancel{s}}{1 \min}) * (\frac{1 mile}{5280 \cancel{ft}}) = ?\]

OpenStudy (anonymous):

write the units, that's the key :-D

OpenStudy (anonymous):

seconds and feet cancel out

OpenStudy (anonymous):

ok so i multiply 240 and 60 then divide by 5280?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

remember your final units

OpenStudy (anonymous):

ok hold on!

OpenStudy (anonymous):

It's what's left

OpenStudy (anonymous):

2.73miles/min?

OpenStudy (anonymous):

Yeah or 30/11 miles per minute

OpenStudy (anonymous):

ok! thank you!